The concentration of lead ions (Pb2+) in a sample of polluted water that also contains nitrate ions (NO3) is determined by adding solid sodium sulfate (Na2SO4) to exactly 500 mL of the water. Calculate the molar concentration of Pb2+ if 0.00230 g of Na2SO4 was needed for the complete precipitation of lead ions as PbSO4.
Answer #1

The reaction between Pb2+ and Na2SO4 is as follows

Pb2+(aq) + Na2SO4 (aq) ------> PbSO4(s) + 2Na+(aq)

stoichiometrically, 1mole of Na2SO4 react with 1mole of Pb2+

moles of Na2SO4 consumed = 0.00230g/142.04g/mol = 1.62×10-5

number of moles of Pb2+ = 1.62×10-5 mol

molarity of Pb2+ =(1.62 × 10-5mol/500ml)×1000ml

Molarity of Pb2+ = 3.24×10-5 M